# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ Appendix B Proof of Technical Lemmas
We now turn our attention to the upper-right 1×($1\times(\ell-1)$ corner of $\sum_{j=1}^{d}P_{j}\mathbb{E}[\xi_{j}\Psi_{j}]P_{j}$. For every unit vector $x\in\mathbb{C}^{\ell}$,
$\left\|\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&\xi_{j}r_{j}\tilde{l}_{-j}^{T}R_{-j}C_{j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{T}R_{-j}C_{j}+\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{T}R_{-j}C_{j}\\ 0&0\end{bmatrix}P_{j}x\right\|_{2}$
$=\left\|\mathbb{E}\begin{pmatrix}\begin{pmatrix}0&\xi_{1}r_{1}\tilde{l}_{-1}^{T}R_{-1}C_{1}-2\xi^{2}_{1}r_{1}^{2}l_{-1}^{T}R_{-1}C_{1}+\xi_{1}(\tilde{q}_{1}-q_{1})l_{-1}^{T}R_{-1}C_{1}\\ \end{pmatrix}P_{1}x\\ \vdots\\ \begin{pmatrix}0&\xi_{d}r_{d}\tilde{l}_{-d}^{T}R_{-d}C_{d}-2\xi^{2}_{d}r_{d}^{2}l_{-d}^{T}R_{-d}C_{d}+\xi_{d}(\tilde{q}_{d}-q_{d})l_{-d}^{T}R_{-d}C_{d}\\ \end{pmatrix}P_{d}x\end{pmatrix}\right\|_{2}$
$\leq\sqrt{\ell}\max_{1\leq j\leq d}\|\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{T}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{T}R_{-j}+\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{T}R_{-j}]\|_{2}.$
@@
$\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{T}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{T}R_{-j}]=\mathbb{E}[\xi_{j}l_{-j}^{T}R_{-j}KR_{-j}-2\xi^{2}_{j}l_{-j}^{T}R_{-j}KR_{-j}l_{-j}l_{-j}^{T}R_{-j}]$
@@
$\max_{1\leq j\leq d}\|\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{T}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{T}R_{-j}]\|\leq\frac{\mathbb{E}[\|l_{-1}\|]\|K\|}{(\Im[z])^{3}}+\frac{2\mathbb{E}[\|l_{-1}\|^{3}]\|K\|}{(\Im[z])^{5}}.$
@@
$\|\mathbb{E}[\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{T}R_{-j}]\|_{2}$ $=\sup_{\|y\|\leq 1}|\mathbb{E}[\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{T}R_{-j}y]|$
$\leq(\Im[z])^{-1}\sup_{\|y\|\leq 1}\sqrt{\mathbb{E}[|q-\tilde{q}|^{2}]\mathbb{E}[|l_{-j}^{T}R_{-j}y|^{2}]}$
@@
Combining everything, we obtain that the upper-right 1×($1\times(\ell-1)$ corner of $\sum_{j=1}^{d}P_{j}\mathbb{E}[\xi_{j}\Psi_{j}]P_{j}$ is bounded, in norm, by
@@
$\underline{u}{\Psi}_{j}=\begin{bmatrix}0&0\\ -2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}-C^{-1}_{j}l_{-j}&C^{-1}_{j}(l_{-j}l_{-j}^{T}-\tilde{l}_{-j}\tilde{l}_{-j}^{T})R_{-j}C_{j}+2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}$
@@
$\mathbb{E}\underline{u}{\Psi}_{j}=\begin{bmatrix}0&0\\ -2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}&2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}.$
@@
$\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$\quad\quad\quad\quad\quad-(1+z+\mathbb{E}q)^{-1}\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&2(q_{j}-\mathbb{E}q_{j})\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$\left\|\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&\begin{subarray}{c}2(q_{j}-\mathbb{E}q_{j})\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{subarray}\end{bmatrix}P_{j}\right\|\leq\frac{2d\|K\|\sqrt{\mathbb{E}[|q-\mathbb{E}q|^{2}]\mathbb{E}[\|l_{-1}\|^{4}]}}{(\Im[z])^{3}}.$

### v4

@@ Appendix B Proof of Technical Lemmas
We now turn our attention to the upper-right $1\times(\ell-1)$ corner of $\sum_{j=1}^{d}P_{j}\mathbb{E}[\xi_{j}\Psi_{j}]P_{j}$. For every unit vector $x\in\mathbb{C}^{\ell}$,
$\left\|\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&\xi_{j}r_{j}\tilde{l}_{-j}^{\top}R_{-j}C_{j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{\top}R_{-j}C_{j}+\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{\top}R_{-j}C_{j}\\ 0&0\end{bmatrix}P_{j}x\right\|_{2}$
$=\left\|\mathbb{E}\begin{pmatrix}\begin{pmatrix}0&\xi_{1}r_{1}\tilde{l}_{-1}^{\top}R_{-1}C_{1}-2\xi^{2}_{1}r_{1}^{2}l_{-1}^{\top}R_{-1}C_{1}+\xi_{1}(\tilde{q}_{1}-q_{1})l_{-1}^{\top}R_{-1}C_{1}\\ \end{pmatrix}P_{1}x\\ \vdots\\ \begin{pmatrix}0&\xi_{d}r_{d}\tilde{l}_{-d}^{\top}R_{-d}C_{d}-2\xi^{2}_{d}r_{d}^{2}l_{-d}^{\top}R_{-d}C_{d}+\xi_{d}(\tilde{q}_{d}-q_{d})l_{-d}^{\top}R_{-d}C_{d}\\ \end{pmatrix}P_{d}x\end{pmatrix}\right\|_{2}$
$\leq\sqrt{\ell}\max_{1\leq j\leq d}\|\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{\top}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{\top}R_{-j}+\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{\top}R_{-j}]\|_{2}.$
@@
$\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{\top}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{\top}R_{-j}]=\mathbb{E}[\xi_{j}l_{-j}^{\top}R_{-j}KR_{-j}-2\xi^{2}_{j}l_{-j}^{\top}R_{-j}KR_{-j}l_{-j}l_{-j}^{\top}R_{-j}]$
@@
$\max_{1\leq j\leq d}\|\mathbb{E}[\xi_{j}r_{j}\tilde{l}_{-j}^{\top}R_{-j}-2\xi^{2}_{j}r_{j}^{2}l_{-j}^{\top}R_{-j}]\|\leq\frac{\mathbb{E}[\|l_{-1}\|]\|K\|}{(\Im[z])^{3}}+\frac{2\mathbb{E}[\|l_{-1}\|^{3}]\|K\|}{(\Im[z])^{5}}.$
@@
$\|\mathbb{E}[\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{\top}R_{-j}]\|_{2}$ $=\sup_{\|y\|\leq 1}|\mathbb{E}[\xi_{j}(\tilde{q}_{j}-q_{j})l_{-j}^{\top}R_{-j}y]|$
$\leq(\Im[z])^{-1}\sup_{\|y\|\leq 1}\sqrt{\mathbb{E}[|q-\tilde{q}|^{2}]\mathbb{E}[|l_{-j}^{\top}R_{-j}y|^{2}]}$
@@
Combining everything, we obtain that the upper-right $1\times(\ell-1)$ corner of $\sum_{j=1}^{d}P_{j}\mathbb{E}[\xi_{j}\Psi_{j}]P_{j}$ is bounded, in norm, by
@@
$\underline{u}{\Psi}_{j}=\begin{bmatrix}0&0\\ -2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}-C^{-1}_{j}l_{-j}&C^{-1}_{j}(l_{-j}l_{-j}^{\top}-\tilde{l}_{-j}\tilde{l}_{-j}^{\top})R_{-j}C_{j}+2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}$
@@
$\mathbb{E}\underline{u}{\Psi}_{j}=\begin{bmatrix}0&0\\ -2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}&2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}$.
@@
$\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&2\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$\quad\quad\quad\quad\quad-(1+z+\mathbb{E}q)^{-1}\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&2(q_{j}-\mathbb{E}q_{j})\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$\left\|\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}0&0\\ 0&\begin{subarray}{c}2(q_{j}-\mathbb{E}q_{j})\xi_{j}C^{-1}_{j}KR_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{subarray}\end{bmatrix}P_{j}\right\|\leq\frac{2d\|K\|\sqrt{\mathbb{E}[|q-\mathbb{E}q|^{2}]\mathbb{E}[\|l_{-1}\|^{4}]}}{(\Im[z])^{3}}.$